OpenStudy (anonymous):

How do I express the answer with positive exponents for this question, (2m^2/n^3)^-2?

OpenStudy (anonymous):

(2m^2/n^3)^-2 1 = ------------ (2m^2/n^3)^2 =[(n^3)/2m^2]^2 answer

OpenStudy (anonymous):

or just reverse the denominator and the numerator

OpenStudy (anonymous):

=n^6/(4m^4)=0.25n^6 /m^4

OpenStudy (anonymous):

any question milliex? have fun now

OpenStudy (anonymous):

thank you so much!